# Intermediate models and Kinna–Wagner Principles

### v2

@@ 1. Introduction
The Kinna–Wagner Conjecture can be understood as a generalisation of the intermediate model theorem, which states that if $V\subseteq M\subseteq V[G]$ are models of $\mathsf{ZFC}$, where $G$ is a $V$-generic filter, then $M$ is a generic extension of $V$ and $V[G]$ is a generic extension of $M$. This is not true if $M$ does not satisfy the Axiom of Choice, as generic extensions preserve the Axiom of Choice. But a natural question is to ask about symmetric extensions instead.
Grigorieff [2] studied intermediate models of $\mathsf{ZF}$, mainly in the form of symmetric extensions, even without the assumption of $\mathsf{AC}$ in $V$ or $V[G]$. The work on the Bristol model in [7, 9] showed that if $c$ is a Cohen real over $L$, then the structure of intermediate models is far more complex than previously imagined. Not every intermediate model has the form $L(x)$ for any set $x$, and even those that do, are not necessarily symmetric extension given by the Cohen real itself. These results were recently extended by Hayut and Shani [3], where they show that given any model of $\mathsf{ZF}$, $V$, if $c$ is a Cohen real over $V$, then there is an intermediate $V\subseteq M\subseteq V[c]$ such that $M\neq V(x)$ for all $x$. Hayut and Shani show that in these intermediate models all Kinna–Wagner Principles fail. We show in the final section of this work that more can be said.
@@ 4.1. Kinna–Wagner in the forcing multiverses
Suppose that $V[G]\models\mathsf{KWP}_{\alpha}^{*}$, where $G\subseteq\mathbb{P}\subseteq\mathcal{P}^{\beta}(\mathrm{Ord})$ is $V$-generic, then $V\models\mathsf{KWP}_{\beta+\alpha}^{*}$.
@@
First, note that in $V$, any $a\in\mathcal{P}^{\alpha}(\mathbb{P}\times\mathrm{Ord})$ codes, uniformly, a $\mathbb{P}$-name, $\tau(a)$, defined recursively by $\tau(\langle p,\xi\rangle)=\check{\xi},\ \tau(a)=\{\langle p,\tau(b)\rangle\mid\langle p,b\rangle\in a\}$. Moreover, for every $z\in\mathcal{P}^{\alpha}(\mathrm{Ord})^{V[G]}$, there is some $a\in\mathcal{P}^{\alpha}(\mathbb{P}\times\mathrm{Ord})$ such that $\tau(a)^{G}=z$. And, of course, there is a surjection from $\mathcal{P}^{\beta+\alpha}(\mathrm{Ord})\to\mathbb{P}\times\mathcal{P}^{\alpha}(\mathbb{P}\times\mathrm{Ord})$.
Now letting $x\in V$ be an arbitrary set, there is a $\mathbb{P}$-name $\dot{f}\in V$ and $p\in\mathbb{P}$ such that $p\mathrel{\Vdash}``\dot{f}\colon\mathcal{P}^{\alpha}(\mathrm{Ord})^{\bullet}\to\check{x}$ is a surjection”. Consider the function $g(p,a)=y$ if and only if $p\mathrel{\Vdash}\dot{f}(\tau(a))=\check{y}$, then $g$ is a surjection from a subset of ×$\mathbb{P}\times\mathcal{P}^{\alpha}(\mathrm{Ord})$ onto $x$. Composing with the surjection, we get a surjection from $\mathcal{P}^{\beta+\alpha}\to x$. ∎
@@ 4.2. Proof of the Kinna–Wagner Conjecture
It is enough to prove this in the case where $\mathbb{Q}$ is a complete Boolean algebra, since every forcing notion embeds densely into its Boolean completion. Let $\dot{x}$ be a $\mathbb{Q}$-name such that $\dot{x}^{H}=x$, and let $\mathbb{Q}_{\dot{x}}$ be the complete subalgebra generated by the conditions of the form $[\![\check{z}\in\dot{x}]\!]$ for $z\in V$. Then $W[H\cap\mathbb{Q}_{\dot{x}}]=W[x]$, see the proof of [5, Corollary 15.42]. [^5] The proof in Jech is in the context of $\mathsf{ZFC}$, but the only key point is that $A\subseteq\kappa$ is a subset of $V$.
@@
First let $\beta$ be the least such that $V\models\mathsf{KWP}_{\beta}^{*}$. Next, for each $z\subseteq\mathcal{P}^{\beta+\alpha}(\eta_{\alpha})$, there is a complete Boolean algebra $\mathbb{Q}\in V(z)$ and a $V(z)$-generic filter, $H$, such that $V(z)[H]=V[G]$. By subsection 3.2, $V(z)\models\mathsf{KWP}_{\beta+\alpha}^{*}$, and therefore there is an ordinal $\eta$ and a surjection from $\mathcal{P}^{\beta+\alpha}(\eta)^{V(z)}$ onto $\mathbb{Q}$ in $V(z)$, we let $\eta^{z}$ be the least such ordinal. Then, let $\eta^{\prime}_{\alpha+1}=\sup\{\eta^{z}\mid z\subseteq\mathcal{P}^{\beta+\alpha}(\eta^{\prime}_{\alpha})\}$.
Let $M$ be an intermediate model of $\mathsf{KWP}^{*}_{\alpha}$. We claim that $M=V(z)$ for some $z\subseteq\mathcal{P}^{\beta+\alpha}(\eta_{\alpha+1})^{M}$. Let $X_{\delta}$ denote $\mathcal{P}^{\delta}(\mathrm{Ord})^{M}$. First, note that $V(X_{\delta})\models\mathsf{KWP}_{\beta+\delta}$, and since $M\models\mathsf{KWP}^{*}_{\alpha}$, $M=L(X_{\alpha+1})=V(X_{\alpha+1})$. We claim that $V(X_{\delta})=V(z)$ for some $z\subseteq\mathcal{P}^{\beta+\delta}(\eta_{\delta})$. We prove this claim by induction on $\delta$. For $\delta=0$, this is trivial, $V(X_{0})=V$ so we can take $z=\varnothing$. If $\delta$ is a limit, by the continuity of $X_{\delta}$, it follows that for $\gamma<\delta$, since $V(X_{\gamma})=V(z_{\gamma})$ for some $z_{\gamma}$, then $V(X_{\gamma})=V(\mathcal{P}^{\beta+\gamma+1}(\eta_{\gamma})^{M})$. Therefore, $V(X_{\delta})=V(\mathcal{P}^{\beta+\delta}(\eta_{\delta})^{M})$ as wanted.
It remains to prove this in the successor case for $\delta+1$. Note that $V(X_{\delta})=V(z)$ for some $z\subseteq\mathcal{P}^{\beta+\delta}(\eta_{\delta}^{\prime})^{M}$, therefore we can fix some $\mathbb{Q}\in V(z)$ such that for some $V(z)$-generic filter, $H$, $V(z)[H]=V[G]$. By subsection 4.2, for every $x\in X_{\delta+1}$, since $x\subseteq V(z)$, there is some $y\subseteq\mathbb{Q}$ such that $V(z,x)=V(z,y)$. Therefore,
$V(X_{\delta+1})\subseteq V(\mathcal{P}(z)^{M})\subseteq V(\mathcal{P}^{\beta+\delta+1}(\eta_{\delta}^{\prime})^{M})\subseteq V(\mathcal{P}^{\beta+\delta+1}(\eta_{\delta+1}^{\prime})^{M}).$
@@ 5. Generalisation and applications
Moreover, it is to verify that $x_{\delta}\times\mathbb{P}\times\mathcal{P}^{\delta}(\mathbb{P}^{<\omega})^{V}$ is in bijection with $x_{\delta}$, so we get a function from $x_{\delta}$ onto $\mathbb{Q}$ whose image is dense. By subsection 4.2 we have that $V(x_{\delta},x)=V(x_{\delta},y)$ for some $y\subseteq x_{\delta}$. Therefore, $V(x_{\delta+1})=X(X_{\delta+1})$ as wanted. ∎
@@
The equivalence between (1) and (2) had been established by Grigorieff [2]. The implication from (1) to (4) follows directly from the definitions of $V(x)$ and $\mathsf{SVC}_{V}$. The proof that (4) implies (3) has a similar proof to subsection 3.2. Finally, the proof of Theorem 4.2 translates directly [^6] Mutantis mutandi. to this case, establishing the implication from (3) to (1) ∎
@@ Acknowledgements
The authors would like to thank Elliot Glazer for suggesting the concept of $\mathsf{KWP}^{*}$. The authors were supported by a UKRI Future Leaders Fellowship [MR/T021705/2]. No data are associated with this article. For the purpose of open access, the authors have applied a CC-BY licence to any Author Accepted Manuscript version arising from this submission.

### v3

@@ 1. Introduction
The Kinna–Wagner Conjecture can be understood as a generalisation of the intermediate model theorem, which states that if $V\subseteq M\subseteq V[G]$ are models of $\mathsf{ZFC}$, where $G$ is a $V$-generic filter, then $M$ is a generic extension of $V$ and $V[G]$ is a generic extension of $M$. This is not true if $M$ does not satisfy the Axiom of Choice, as generic extensions preserve the Axiom of Choice. But a natural question is whether $M$ must be a symmetric extension of $V$, at least in the presence of the Axiom of Choice, or some Kinna–Wagner Principle.
Grigorieff [2] studied intermediate models of $\mathsf{ZF}$, mainly in the form of symmetric extensions, even without the assumption of $\mathsf{AC}$ in $V$ or $V[G]$. The work on the Bristol model in [7, 9] showed that if $c$ is a Cohen real over $L$, then the structure of intermediate models is far more complex than previously imagined. Not every intermediate model has the form $L(x)$ for any set $x$, and even those that do, are not necessarily symmetric extension given by the Cohen real itself. These results were recently extended by Hayut and Shani [3], where they show that given any model of $\mathsf{ZF}$, $V$, if $c$ is a Cohen real over $V$, then there is an intermediate $V\subseteq M\subseteq V[c]$ such that $M\neq V(x)$ for all $x$. Hayut and Shani show that in these intermediate models all Kinna–Wagner Principles fail.
We show in the final section of this work that more can be said. Theorem 4.2 shows that if $V\subseteq M\subseteq V[G]$ are models of $\mathsf{ZF}$, with $G$ a $V$-generic filter, and some Kinna–Wagner Principle holds in $V$, then $M$ has the form $V(x)$ for some set $x$ if and only if it satisfies some Kinna–Wagner Principle, and that the set $x$ is uniformly defined from the Kinna–Wagner Principle. The notion of a Kinna–Wagner Principle is generalised to allow different “starting classes” than the ordinals, and consequently we obtain Theorem 5.2.
@@ 4.1. Kinna–Wagner in the forcing multiverses
If $V[G]\models\mathsf{KWP}_{\alpha}^{*}$, where $G\subseteq\mathbb{P}\subseteq\mathcal{P}^{\beta}(\mathrm{Ord})$ is $V$-generic, then $V\models\mathsf{KWP}_{\beta+\alpha}^{*}$.
@@
We define a sequence of classes of $\mathbb{P}$-names, $S_{\delta}$, by recursion:
  1. (1)
    $S_{0}=\{\check{\xi}\mid\xi\in\mathrm{Ord}\}$,
  2. (2)
    $S_{\delta+1}=\mathcal{P}(\mathbb{P}\times S_{\delta})$,
  3. (3)
    $S_{\delta}=\bigcup_{\gamma<\delta}S_{\gamma}$ for a limit ordinal $\delta$.
If $a\in\mathcal{P}^{\delta}(\mathrm{Ord})^{V[G]}$, then there is some $\dot{a}\in S_{\delta}$ such that $\dot{a}^{G}=a$. Moreover, in $V$ there is a definable surjection $\tau_{\delta}\colon\mathcal{P}^{\beta+\delta}(\mathrm{Ord})\to S_{\delta}$, since $\mathbb{P}\subseteq\mathcal{P}^{\beta}(\mathrm{Ord})$.
Now letting $x\in V$ be an arbitrary set, there are $\mathbb{P}$-names $\dot{a}\in S_{\alpha+1}$ and $\dot{f}$, and a condition $p\in G$ such that $p\mathrel{\Vdash}``\dot{f}\colon\dot{a}\to\check{x}$ is a surjection”. Let $g$ be the function $g(q,s)=y$ if and only if $q\mathrel{\Vdash}\dot{f}(\tau_{\alpha}(s))=\check{y}$, then $g$ is a surjection from a subset of ×$\mathbb{P}\times\mathcal{P}^{\beta+\alpha}(\mathrm{Ord})$ onto $x$. So $V\models\mathsf{KWP}_{\beta+\alpha}^{*}$, as wanted. ∎
@@ 4.2. Proof of the Kinna–Wagner Conjecture
It is enough to prove this in the case where $\mathbb{Q}$ is a complete Boolean algebra, since every forcing notion embeds densely into its Boolean completion. Let $\dot{x}$ be a $\mathbb{Q}$-name such that $\dot{x}^{H}=x$, and let $\mathbb{Q}_{\dot{x}}$ be the complete subalgebra generated by the conditions of the form $[\![\check{z}\in\dot{x}]\!]$ for $z\in V$. Then $W[H\cap\mathbb{Q}_{\dot{x}}]=W[x]$, see the proof of [5, Corollary 15.42]. [^5] The proof in Jech is in the context of $\mathsf{ZFC}$, but the Axiom of Choice is not used in the proof beyond the fact that $\mathsf{KWP}_{0}$ and the Balcar–Vopěnka theorem reduce all cases to sets of ordinals and that sets of ordinals are subsets of $V$.
@@
First let $\beta$ be the least such that $V\models\mathsf{KWP}_{\beta}^{*}$. Next, for each $z\subseteq\mathcal{P}^{\beta+\alpha}(\eta_{\alpha}^{\prime})$, there is a complete Boolean algebra $\mathbb{Q}\in V(z)$ and a $V(z)$-generic filter, $H$, such that $V(z)[H]=V[G]$. By subsection 3.2, $V(z)\models\mathsf{KWP}_{\beta+\alpha}^{*}$, and therefore there is an ordinal $\eta$ and a surjection from $\mathcal{P}^{\beta+\alpha}(\eta)^{V(z)}$ onto $\mathbb{Q}$ in $V(z)$, we let $\eta^{z}$ be the least such ordinal. Then, let $\eta^{\prime}_{\alpha+1}=\sup\{\eta^{z}\mid z\subseteq\mathcal{P}^{\beta+\alpha}(\eta^{\prime}_{\alpha})\}$.
Let $M$ be an intermediate model of $\mathsf{KWP}^{*}_{\alpha}$. We claim that $M=V(z)$ for some $z\subseteq\mathcal{P}^{\beta+\alpha}(\eta^{\prime}_{\alpha+1})^{M}$. Let $X_{\delta}$ denote $\mathcal{P}^{\delta}(\mathrm{Ord})^{M}$. First, note that $V(X_{\delta})\models\mathsf{KWP}_{\beta+\delta}$, and since $M\models\mathsf{KWP}^{*}_{\alpha}$, $M=L(X_{\alpha+1})=V(X_{\alpha+1})$. We claim that $V(X_{\delta})=V(z)$ for some $z\subseteq\mathcal{P}^{\beta+\delta}(\eta^{\prime}_{\delta})$. We prove this claim by induction on $\delta$. For $\delta=0$, this is trivial, $V(X_{0})=V$ so we can take $z=\varnothing$. If $\delta$ is a limit, by the continuity of $X_{\delta}$, it follows that for $\gamma<\delta$, since $V(X_{\gamma})=V(z_{\gamma})$ for some $z_{\gamma}$, then $V(X_{\gamma})=V(\mathcal{P}^{\beta+\gamma+1}(\eta^{\prime}_{\gamma})^{M})$. Therefore, $V(X_{\delta})=V(\mathcal{P}^{\beta+\delta}(\eta^{\prime}_{\delta})^{M})$ as wanted.
It remains to prove this in the successor case for $\delta+1$. Note that $V(X_{\delta})=V(z)$ for some $z\subseteq\mathcal{P}^{\beta+\delta}(\eta_{\delta}^{\prime})^{M}$, therefore we can fix some $\mathbb{Q}\in V(z)$ such that for some $V(z)$-generic filter, $H$, $V(z)[H]=V[G]$, by the choice of $\eta_{\delta+1}^{\prime}$ there is a surjection from $\mathcal{P}^{\beta+\delta}(\eta_{\delta+1}^{\prime})^{V(z)}$ onto $\mathbb{Q}$ in $V(z)$. By subsection 4.2, for every $x\in X_{\delta+1}$, since $x\subseteq V(z)$, there is some $y\subseteq\mathcal{P}^{\beta+\delta}(\eta_{\delta+1}^{\prime})$ such that $V(z,x)=V(z,y)$. Therefore,
$V(X_{\delta+1})\subseteq V(\mathcal{P}^{\beta+\delta+1}(\eta_{\delta+1}^{\prime})^{M}).$
Therefore, $V(X_{\delta+1})=V(\{y\subseteq\mathcal{P}^{\beta+\delta}(\eta_{\delta+1}^{\prime})^{M}\mid V(z^{\prime})\subseteq V(X_{\delta+1})\})$.
@@ 5. Generalisation and applications
Moreover, it is to verify that $x_{\delta}\times\mathbb{P}\times\mathcal{P}^{\delta}(\mathbb{P}^{<\omega})^{V}$ is in bijection with $x_{\delta}$, so we get a function from $x_{\delta}$ onto $\mathbb{Q}$ whose image is dense. By subsection 4.2 we have that $V(x_{\delta},x)=V(x_{\delta},y)$ for some $y\subseteq x_{\delta}$. Therefore, $V(x_{\delta+1})=V(X_{\delta+1})$. ∎
@@
The equivalence between (1) and (2) had been established by Grigorieff [2]. The implication from (1) to (4) follows directly from the definitions of $V(x)$ and $\mathsf{SVC}_{V}$. The proof that (4) implies (3) has a similar proof to subsection 3.2. Finally, the proof of Theorem 4.2 translates directly [^6] Mutantis mutandi. to this case, establishing the implication from (3) to (1). ∎
@@ Acknowledgements
The authors would like to thank Elliot Glazer for suggesting the concept of $\mathsf{KWP}^{*}$, as well as the anonymous referee for their helpful suggestions. The authors were supported by a UKRI Future Leaders Fellowship [MR/T021705/2]. No data are associated with this article. For the purpose of open access, the authors have applied a CC-BY licence to any Author Accepted Manuscript version arising from this submission.